# Statistical reduced order modelling for the parametric Helmholtz equation

### v1

@@ 4.4.1 Hyperparameter Learning
To learn the posterior hyperparameters from data, the marginal likelihood has to be derived for the statROM case, as well. For a single observation there holds for the statROM data model
@@
with $n_{y}$ the number of sensors $\bm{y}_{i}$ and
@@
The marginal likelihood is minimized using a gradient based optimizer. In this case, L-BFGS-B [40] is used. It is visible in (64) that the misspecified ROM prior mean at the sensor locations $\rho\bm{P}\bm{\mu}_{\tilde{\bm{u}}}$ is corrected by the mean error estimate $\rho\bm{P}\bm{\mu}_{\bm{d}_{r}}$.
@@
By a slight abuse of notation, we denote with $\tilde{\bm{u}}_{Y_{d_{r}}}$ the entire predictive field. It is visible that the posterior mean and covariance are corrected by the model error and ROM error terms. Following the reasoning in [4], the predictive observation density now tends, in contrast to the plain posterior, to $\bm{y}$ for the mean and to $2\bm{C}_{d}+\bm{C}_{d_{r}}+\bm{C}_{e}$ for the covariance, hence to the data generating process.
@@
5:     $\bm{M}^{(i)},\,\bm{S}^{(i)},\bm{D}^{(i)}=\mathrm{assembleSystem}(\bm{\xi}^{(i)})$triangleright FEM system assembly
6:     $\bm{V}^{(i)},\bm{L}^{(i)},\bm{U}^{(i)}=\mathrm{AORA}(\bm{M}^{(i)},\bm{S}^{(i)},\bm{D}^{(i)},\bm{f}^{(i)},\bar{\omega},m)$triangleright Primal ROM basis
@@
8:         $\bm{V}_{\mathrm{adj},\ell}^{(i)}=\mathrm{AORA}((\bm{M}^{(i)})^{*},(\bm{S}^{(i)})^{*},(\bm{D}^{(i)})^{*},\bm{p}_{\ell},\bar{\omega},m,\bm{L}^{(i)},\bm{U}^{(i)})$triangleright Adjoint ROM basis
@@
1:$\omega\in[\underline{\omega},\overline{\omega}]$, $\bar{\omega}$, $m$, $c$, $\bm{V}^{n_{\mathrm{qmc}}}$, $\bm{V}_{\mathrm{adj}}^{n_{\mathrm{qmc}}\times n_{\mathrm{adj}}}$
2:$\omega$ non-resonant $\triangleright$ (For an undampened system)
3:$k=\omega/c$
@@
5:     $\bm{M}^{(i)},\,\bm{S}^{(i)},\bm{D}^{(i)}=\mathrm{assembleSystem}(\bm{\xi}^{(i)})$triangleright FEM system assembly
6:     $\bm{A}^{(i)}(\omega)=(-k^{2}\bm{M}^{(i)}+k\bm{D}^{(i)}+\bm{S}^{(i)})$
@@
8:     $\bm{u}_{r}^{(i)}=(\bm{A}_{r}^{(i)})^{-1}(\bm{f}_{r}^{(i)}+\bm{f}_{g,r}^{(i)})$
9:     $\bm{d}_{r}^{(i)}=\mathrm{adjEst}(\bm{V}_{\mathrm{adj}}^{(i)},\bm{A}^{(i)},\bm{f}_{r}^{(i)}+\bm{f}_{g,r}^{(i)},\bm{u}_{r}^{(i)})$triangleright Adjoint error estimation (47)
@@
11:$\bm{U}_{r}=[\bm{u}_{r}^{(1)},\dots,\bm{u}_{r}^{(I)}]$
12:$\bm{D}_{r}=[\bm{d}_{r}^{(1)},\dots,\bm{d}_{r}^{(I)}]$
13:$\tilde{\bm{u}}\sim\mathcal{N}(\mathrm{mean}(\bm{V}\bm{U}_{r}),\mathrm{cov}(\bm{V}\bm{U}_{r}))$triangleright ROM prior (32)
14:$\bm{d}_{r}\sim\mathcal{N}(\mathrm{mean}(\bm{D}_{r}),\mathrm{cov}(\bm{D}_{r}))$triangleright define the ROM error, GP training
15:$\tilde{\bm{u}},\bm{d}_{r},\bm{Y}$, $\bm{P}$triangleright Measurement data and sensor locations
16:$\tilde{\bm{u}}_{Y_{d_{r}}}=\mathrm{statROM}(\tilde{\bm{u}},\bm{d}_{r},\bm{Y},\bm{P})$triangleright Compute posterior state (61)
@@ 5 Numerical Examples
We consider two numerical examples to illustrate the new methods. The first one is a 1D example of the Helmholtz equation with random forcing and a random heterogeneous material coefficient. It is used to show the convergence properties of the proposed statROM approach compared to the classical approach without a ROM error estimator. The second example is more involved and deals with solving the Helmholtz equation in a 2D scattering scenario with a misspecified random forcing. For both settings, we compare the predictive observation density derived from the classical data model (38) and from the proposed statROM data model (67). In both cases, the prior state is computed using the ROM. We denote the predictive state for the classical approach with $\tilde{\bm{u}}_{Y}$ and for the proposed statROM approach with $\tilde{\bm{u}}_{Y_{d_{r}}}$.
@@ 5.1 One-dimensional problem
To illustrate the method and to show the convergence properties, we begin by investigating the Helmholtz equation in 1D. We consider an inhomogeneous Neumann boundary condition at $x_{n}=0$ and no impedance boundary condition, hence, the problem is real-valued. We choose $f_{g}\sim\mathcal{N}(\frac{\pi^{2}}{50},0.02^{2})$. At $x_{n}=1$, a homogeneous Neumann boundary condition, i.e. a sound hard reflective boundary, is applied. The material parameter random field is parameterized with $\sigma_{\kappa}=\sqrt{5e-2},\;l_{\kappa}=0.3$ for an exponential kernel function. The FE mesh consists of $100$ elements and linear Ansatz functions are employed. We first investigate the convergence of the ROM prior itself. Then we condition with the standard statFEM error model to obtain a reference, followed by a statROM update. Finally, we investigate the convergence of the predictive posterior solution towards the data generating process.
In Figure 1, the 1D problem setup is exemplary shown for $\omega/2\pi=460$Hz. We generate synthetic data, by adding noisy perturbations to a full order reference $\bm{\mu}_{\bm{u};\mathrm{ref}}(\omega)$ with $N_{\mathrm{ref}}=1000$ DOFs. The data are collected on $n_{y}=11$ equally spaced sensor locations throughout the domain. Per sensor, $n_{o}=20$ observations are taken. The FEM prior $\bm{u}$ is the solution of an FE simulation, obtained with the same setting as $\bm{\mu}_{\bm{u};\mathrm{ref}}(\omega)$, with $N=101$ DOFs and QMC sampling. The model error, hence, is only due to discretization. The ROM prior $\tilde{\bm{u}}$, which is shown here for a moment matching ROM with $m=5$ moments matched around $\bar{\omega}/2\pi=100$Hz, neither follows the data nor the FEM prior. Hence, there clearly is a significant ROM approximation error. This error is supposed to decrease with more matched moments $m$. In the next Section, the convergence of the ROM prior towards the FEM prior with increasing $m$ is analyzed.
Figure 1: Full order and ROM priors for the 1D problem. As the ROM is constructed with only $m=5$ moments matched around the expansion frequency, the approximation error is large. Observation data is collected on a finely discretized reference FE solution on $11$ equally spaced sensor locations. The shaded areas are the $2\sigma$ bands for the respective prior.
@@ 5.1.1 ROM convergence
Figure 2 shows the convergence of the moment matching ROM. The expansion is centered around a single expansion frequency $\bar{\omega}/2\pi=100\mathrm{Hz}$. The ROM solution $\tilde{\bm{u}}(\omega)$ and the full order solution $\bm{u}(\omega)$ are evaluated at $\omega/2\pi=460\mathrm{Hz}$. It is visible that the ROM converges, in this case, to machine precision at $m=15$ moments chosen for the expansion.
Figure 2: The approximation error of the ROM mean quickly converges to a constant quantity at $m=15$ matched moments.
Choosing a fixed number of moments in the expansion and calculating the error norm for a set of frequencies yields the result visible in Figure 3. As the expansion is centered around $\bar{\omega}/2\pi=100\mathrm{Hz}$, the error at that frequency is very low, regardless of how many moments are used in the expansion. Clearly, the more moments are used, the lower the error also for frequencies further away from $\bar{\omega}$.
Figure 3: For higher numbers of matched moments $m$ around $\bar{\omega}/2\pi=100$Hz, the ROM yields smaller errors throughout the frequency range. With $m=15$, the error is similarly low throughout the frequency range.
@@
Throughout this section, the performance of the proposed method is compared to a reference, where the standard (statFEM) posterior update is performed on a ROM prior instead of a FEM prior. The prior is conditioned on noisy data drawn from a finely discretized ($1000$ elements) FEMsolution of the problem. For the statROM posterior update, the ROM error estimate is used to construct the predictive obervation density, which corrects the biased posterior with the estimated error. In the next paragraphs, the error between the data generating process and the results of the different approaches is evaluated for different ROM fidelities to show the convergence properties.
@@
The adjoint error estimator is evaluated at a set of training points throughout the domain. For each of the points, a ROM projection matrix has been constructed. Figure 4 shows the convergence of the ROM error estimate with $12$ training points. It is visible that for ROMs which have been constructed with a rather low number of moments, also the ROM error estimate is not as accurate as for higher fidelity ROMs. This is an interesting fact which sets the bounds in which the proposed method is applicable: For very poor quality ROMs, e.g. as shown in Figure 1, the error estimator will also be of rather low quality and no improvement over the standard method will be obtained. On the other hand, if the ROM has a very accuarcy, the ROM error estimator is highly accurate, but the error is so small that it doesn’t need to be quantified. The method is most relevant, as will be shown in the next chapters, for ROMs of medium quality. There, the estimation of the ROM error is feasible with a good accuarcy and the errors are still large enough so that an error correction improves the posterior update.
Figure 4: Convergence of the ROM error estimator at the training points with increasing number of matched moments. A relative error of $0.1\%$ is reached at about $15$ matched moments.
Figures 5 and 6 demonstrate the effect of using more training points for the error GP as well as a higher fidelity ROM. The ROM error decreases and the quality of the estimator increases. The convergence of the estimator at the test points (as opposed to the training points which have been covered in Fig. 4) depends on the chosen GP configuration: Besides the number of training points, also other factors such as the choice of kernel function and hyperparameters have an effect on the absolute error and convergence. The reader is referred to [5] for details on how a GP regression can be optimized. In the following chapters, a Matérn $5/2$ kernel function and frequency dependent hyperparameters are used, i.e. the length scale
@@
Figure 5: For the example from before with $\bar{\omega}/2\pi=100$Hz, $m=5$, the ROM error is approximated with an adjoint error indicator which is evaluated at $14$ training points. A GP is trained to give an approximation of the ROM error in the whole field.
Figure 6: Increasing the number of test points to $22$ and choosing a $6$ instead of $5$ matched moments leads to a better approximation of the ROM error. The amplitude of the error naturally decreases because of the larger ROM basis.
@@
In Figure 7 the convergence of different ROM predictive observation densities is displayed as a function of the chosen basis size. The baseline error is determined by the full order statFEM predictive density, which still has a certain error compared to the data generating process. This error depends on remaining discretization error (data is generated on a very fine grid), the finite number of observations per sensor, the defined sensor noise, the choice of norm and also the small diagonal quantity added to the prior covariance matrices for numerical stability. Plotted in bold lines are the errors of the predictive densities constructed with the proposed adjoint ROM error estimator. For comparison, the convergence of the standard approach, i.e without any ROM error estimation is shown, too. It is visible that a simple version of the adjoint estimator, i.e. with $12$ training points, already delivers a faster convergence than the classical approach.
@@
Figure 7: The expected value of the error for different sizes of the ROM basis. The classical (statFEM) approach applied on the ROM prior yields a slower convergence than the statROM approach with the simplest adjoint error estimator. Increasing the fidelity of the error estimator yields even faster convergence.
The corresponding predictive states for $m=5$ are visible in Figure 8. The ROM error estimate was computed with $12$ training points evenly spaced throughout the domain. In the result for the standard approach, i.e. the yellow curve $\tilde{\bm{u}}_{Y}$, the ROM prior is scaled with $\rho$ to account for the data as good as possible. The estimated model error is quite large as the ROM approximation error is part of it. Hence, the resulting variance is also relatiely large compared to the prior and the full order predictive posterior. The statROM approach with the adjoint ROM error estimator, i.e. the purple line $\tilde{\bm{u}}_{Y;d_{r}}$, results in a smaller model error, hence a much smaller predictive posterior variance, and a mean much closer to the data generating process.
@@
Figure 8: A comparison of the resulting posterior solutions from the ROM prior, the ROM prior with the error estimator and the full order reference prior. The model error is overestimated for low fidelity ROMs because the ROM error is considered model error as well. To illustrate the principles, a bad ROM representation with basis size $m=5$ has been chosen on purpose. It is visible that the estimated ROM error is used as a correction bias for the data (red dots). The data corrected by the error estimator lie much closer to the ROM prior. Hence, the resulting model error for the proposed approach is much smaller than for the classical one, in which the ROM prior is conditioned on uncorrected data. Clearly, using the estimated ROM error to correct the posterior yields a result close to the data generating process even though the ROM prior is used.
@@
It is visible that the proposed statROM approach yields more accurate results. The key insight of the results in Figure 8 is the data correction induced by the statROM update. Clearly, the yellow curve (statFEM) is the ROM prior scaled by $\rho$ so that it passes through the measurement data. Because of the ROM-corrected data model, here the resulting marginal likelihood (64) introduces a bias as big as the ROM error at the data points: the hyperparameters are estimated on order to comply with the corrected data, i.e. data close to the ROM prior. Therefore, the estimated model error GP ($\bm{d}$) of $\tilde{\bm{u}}_{Y_{d_{r}}}$ exhibits a very small $\sigma_{d}$: The ROM-prior lies close to the corrected data. The variance for the corrected posterior, as visible in Fig. 8, is rather small.
@@
On the other hand, the estimated model error GP $\bm{d}$ of $\tilde{\bm{u}}_{Y}$, i.e. for the uncorrected posterior, exhibits a rather large $\sigma_{d}$: The (ROM)-prior doesn’t lie on the actual measured data. This reflects in a large posterior variance as visible in Fig. 8.

### v2

@@ 4.4.1 Hyperparameter Learning
To learn the posterior hyperparameters from data, the marginal likelihood has to be derived for the statROM case, as well. For a single observation the following holds for the statROM data model:
@@
with $n_{y}$ the number of sensors and
@@
The marginal likelihood is minimized using a gradient based optimizer. In this case, L-BFGS-B [50] is used. Notice that in (67) the misspecified ROM prior mean at the sensor locations $\rho\bm{P}\bm{\mu}_{\tilde{\bm{u}}}$ is corrected by the mean error estimate $\rho\bm{P}\bm{\mu}_{\bm{d}_{r}}$.
@@
By a slight abuse of notation, we denote with $\tilde{\bm{u}}_{Y;{d_{r}}}$ the entire predictive field. One may observe that the posterior mean and covariance are corrected by the model error and ROM error terms. Following the reasoning in [4], the predictive observation density now tends, in contrast to the plain posterior, to $\bm{y}$ for the mean and to $2\bm{C}_{d}+\bm{C}_{d_{r}}+\bm{C}_{e}$ for the covariance, hence to the ground truth.
@@
5:  $\bm{M}^{(i)},\,\bm{S}^{(i)},\bm{D}^{(i)}=\mathrm{assembleSystem}(\bm{\xi}^{(i)})$ $\triangleright$ FEM system assembly
6:  $\bm{V}^{(i)},\bm{L}^{(i)},\bm{U}^{(i)}=\mathrm{AORA}(\bm{M}^{(i)},\bm{S}^{(i)},\bm{D}^{(i)},\bm{F}^{(i)},\bar{\omega},m)$ $\triangleright$ Primal ROM basis
@@
8:   $\bm{V}_{\mathrm{adj},\ell}^{(i)}=\mathrm{AORA}((\bm{M}^{(i)})^{*},(\bm{S}^{(i)})^{*},(\bm{D}^{(i)})^{*},\bm{p}_{\ell},\bar{\omega},m,\bm{L}^{(i)},\bm{U}^{(i)})$ $\triangleright$ Adjoint ROM basis
@@
[…]
@@
5:  $\bm{M}^{(i)},\,\bm{S}^{(i)},\bm{D}^{(i)}=\mathrm{assembleSystem}(\bm{\xi}^{(i)})$ $\triangleright$ FEM system assembly
6:  $\bm{A}^{(i)}(\omega)=(-k^{2}\bm{M}^{(i)}-ik\beta\bm{D}^{(i)}+\bm{S}^{(i)})$
@@
8:  $\bm{u}_{r}^{(i)}=(\bm{A}_{r}^{(i)})^{-1}\bm{F}_{r}^{(i)}$
9:  $\bm{d}_{r}^{(i)}=\mathrm{adjEst}(\bm{V}_{\mathrm{adj}}^{(i)},\bm{A}^{(i)},\bm{F}_{r}^{(i)},\bm{u}_{r}^{(i)})$ $\triangleright$ Adjoint error estimation (50)
@@
[…]
@@ 5 Numerical Examples
We consider two numerical examples to illustrate the new methods. The first one is a 1D example of the Helmholtz equation with random forcing and a random heterogeneous material coefficient. It is used to show the convergence properties of the proposed statROM approach compared to the standard approach without a ROM error estimator. The second example is more involved and deals with solving the Helmholtz equation in a 2D scattering scenario with a misspecified random forcing. For both settings, we compare the predictive observation density derived from the standard data model (41) and from the proposed statROM data model (70). In both cases, the prior state is computed using the ROM. We denote the predictive state for the standard approach with $\tilde{\bm{u}}_{Y}$ and for the proposed statROM approach with $\tilde{\bm{u}}_{Y;{d_{r}}}$.
@@ 5.1 One-dimensional problem
To illustrate the method and to show the convergence properties, we begin by investigating the Helmholtz equation in 1D. We consider an inhomogeneous Neumann boundary condition at $x_{n}=0$ and no impedance boundary condition, hence, the problem is real-valued. We choose $f_{g}\sim\mathcal{N}(\frac{\pi^{2}}{50},0.02^{2})$. At $x_{n}=1$, a homogeneous Neumann boundary condition, i.e. a sound hard reflective boundary, is applied. The material parameter random field is parameterized with $\sigma_{\kappa}=\sqrt{5e-2},\;l_{\kappa}=0.3$ for an exponential kernel function. The FE mesh consists of $100$ elements and linear Ansatz functions are employed. We first investigate the convergence of the ROM prior itself. Then we condition with the standard statFEM error model to obtain a reference, followed by a statROM update. Finally, we investigate the convergence of the predictive posterior solution towards the ground truth.
In Figure 2, the 1D problem setup is exemplary shown for $\omega/2\pi=460$Hz. We generate synthetic data, by adding noisy perturbations to a full order reference $\bm{\mu}_{\bm{u};\mathrm{ref}}(\omega)$ with $N_{\mathrm{ref}}=1000$ DOFs. The data are collected on $n_{y}=11$ equally spaced sensor locations throughout the domain. Per sensor, $n_{o}=20$ observations are taken. The FEM prior $\bm{u}$ is the solution of an FE simulation, obtained with the same setting as $\bm{\mu}_{\bm{u};\mathrm{ref}}(\omega)$, with $N=101$ DOFs and QMC sampling. The model error, hence, is only due to discretization. The ROM prior $\tilde{\bm{u}}$, which is shown here for a moment matching ROM with $m=5$ moments matched around $\bar{\omega}/2\pi=100$Hz, neither follows the data nor the FEM prior. Hence, there clearly is a significant ROM approximation error. This error is supposed to decrease with more matched moments $m$. This is exemplary shown in Figure 3. Notice how a change from $m=5$ to $m=15$ leads to a ROM prior with almost no difference to the full order reference. In the next section, the convergence of the ROM prior towards the FEM prior with increasing $m$ is analysed in more detail.
Figure 2: Full order and ROM priors for the 1D problem. As the ROM is constructed with only $m=5$ moments matched around the expansion frequency, the approximation error is large. Observation data are collected on a finely discretized reference FE solution on $11$ equally spaced sensor locations. At each sensor, multiple noisy measurements (red dots) are recorded.
The shaded areas are the $2\sigma$ bands for the respective prior. The FEM prior mean and the ground truth align because no artificial model error is introduced. The only difference between both solutions is the finer discretisation of the ground truth.
Figure 3: Comparison of the quality of different ROM priors for the 1D problem. With $m=15$ moments matched, the approximation error is negligible compared to $m=5$ moments. The difference between the full order reference mean and the respective ROM prior mean is plotted in gray lines. Notice that for $m=15$ there is no visible deviation from the full order mean, even though a far smaller system of equations was solved. In the $2\sigma$ bands, for $m=15$ only a small error is visible compared to the $m=5$ case.
@@ 5.1.1 ROM convergence
Figure 4 shows the convergence of the moment matching ROM. In the left plot, the expansion is centered around a single expansion frequency $\bar{\omega}/2\pi=100\mathrm{Hz}$. The ROM solution $\tilde{\bm{u}}(\omega)$ and the full order solution $\bm{u}(\omega)$ are evaluated at $\omega/2\pi=460\mathrm{Hz}$. One may observe that the ROM converges, in this case, to machine precision at $m=15$ moments chosen for the expansion.
Choosing a fixed number of moments in the expansion and calculating the error norm for a set of frequencies yields the result plotted on the right side. As the expansion is centered around $\bar{\omega}/2\pi=100\mathrm{Hz}$, the error at that frequency is very low, regardless of how many moments are used in the expansion. Clearly, the more moments are used, the lower the error also for frequencies further away from $\bar{\omega}$.
Figure 4: Left: The approximation error of the ROM mean quickly converges to a constant quantity at $m=15$ matched moments. Therefore, a ROM with $m\geq 16$ can be considered to provide an exact representation of the full order model. For the computation of the error norm, the ROM solution $\tilde{\bm{u}}(\omega)$ and the full order solution $\bm{u}(\omega)$ are evaluated at $\omega/2\pi=460\mathrm{Hz}$ while the expansion frequency is $\omega/2\pi=100\mathrm{Hz}$.
Right: For higher numbers of matched moments $m$ around $\bar{\omega}/2\pi=100$Hz, the ROM yields smaller errors throughout the frequency range. With $m=15$, the error is similarly low throughout the frequency range.
@@
Throughout this section, the performance of the proposed method is compared to a reference, where the standard (statFEM) posterior update is performed on a ROM prior instead of a FEM prior. The prior is conditioned on noisy data drawn from a finely discretized ($1000$ elements) FEM solution of the problem. For the statROM posterior update, the ROM error estimate is used to construct the predictive obervation density, which corrects the biased posterior with the estimated error. In the next paragraphs, the error between the ground truth and the results of the different approaches is evaluated for different ROM fidelities to show the convergence properties.
@@
The adjoint error estimator is evaluated at a set of training points throughout the domain. For each of the points, a ROM projection matrix has been constructed. Figure 5 shows the convergence of the ROM error estimate with $12$ training points. One may observe as the number of matched moments grows, not only the ROM itself but also the corresponding error estimator becomes significantly more accurate. This is an interesting fact which sets the bounds in which the proposed method is applicable: For very poor quality ROMs, e.g. as shown in Figure 2, the error estimator will also be of rather low quality and no improvement over the standard method will be obtained. On the other hand, if the ROM has very high accuracy, the ROM error estimator is highly accurate, but the error is so small that it doesn’t need to be quantified. The method is most relevant, as will be shown in the next chapters, for ROMs of medium quality. There, the estimation of the ROM error is feasible with a good accuarcy and the errors are still large enough so that an error correction improves the posterior update.
Figure 5: Left: Convergence of the ROM error estimator at the training points with increasing number of matched moments. A relative error of $0.1\%$ is reached at about $15$ matched moments.
Right: For the example from before with $\bar{\omega}/2\pi=100$Hz, $m=5$, the ROM error is approximated with an adjoint error indicator which is evaluated at $14$ training points. A GP is trained to give an approximation of the ROM error in the whole field. Notice that while the rough shape of the error is well approximated, there is still a significant difference to the exact ROM error.
Increasing the number of test points to $22$ and choosing $m=6$ instead of $m=5$ matched moments leads to a better approximation of the ROM error. The amplitude of the error naturally decreases because of the larger ROM basis.
Figure 5, right side, demonstrates the effect of using more training points for the error GP as well as a higher fidelity ROM. The ROM error decreases and the quality of the estimator increases. The convergence of the estimator at the test points (as opposed to the training points which have been covered in Fig. 5, left side) depends on the chosen GP configuration: Besides the number of training points, also other factors such as the choice of kernel function and hyperparameters have an effect on the absolute error and convergence. The reader is referred to [5] for details on how a GP regression can be optimized. In the following chapters, a Matérn $5/2$ kernel function and frequency dependent hyperparameters are used, i.e. the length scale
@@
In Figure 6 the convergence of different ROM predictive observation densities is displayed as a function of the chosen basis size. The baseline error is determined by the full order statFEM predictive density, which still has a certain error compared to the ground truth. This error depends on the remaining discretization error (data are generated on a very fine grid), the finite number of observations per sensor, the defined sensor noise, the choice of norm and also the small diagonal quantity added to the prior covariance matrices for numerical stability. Plotted in bold lines are the errors of the predictive densities constructed with the proposed adjoint ROM error estimator. For comparison, the convergence of the standard approach, i.e without any ROM error estimation is shown, too. Notice that a simple version of the adjoint estimator, i.e. with $12$ training points, already delivers faster convergence than statFEM with a ROM prior only, i.e. without the error estimator.
@@
Figure 6: The expected value of the error for different sizes of the ROM basis, i.e. different values of $m$. The standard (statFEM) approach applied to the ROM prior, i.e. $\mu_{\tilde{u}|\bm{Y}}$ in orange, yields a slower convergence than the statROM approach with the simplest adjoint error estimator, i.e. $\mu_{\tilde{u}|\bm{Y};d_{r};12}$ in red. Notice that increasing the fidelity of the error estimator yields even faster convergence, i.e. $\mu_{\tilde{u}|\bm{Y};d_{r};22}$ in green.
@@
Figure 7: A comparison of the resulting posterior solutions from the ROM prior $\tilde{u}_{|Y}$ in yellow, the ROM prior with the error estimator $\tilde{u}_{|Y;d_{r}}$ in purple (the proposed statROM approach) and the full order reference prior $u_{|Y}$ in blue. Notice that for the standard approach, i.e. $\tilde{u}_{|Y}$, the model error is overestimated for low fidelity ROMs because the ROM error is considered model error as well. This results in a large posterior standard deviaton. To illustrate the principles, a bad ROM representation with basis size $m=5$ has been chosen on purpose. One may observe that the estimated ROM error is used as a correction bias for the data (red dots). The data corrected by the error estimator lie much closer to the ROM prior. Hence, the resulting model error for the proposed approach is much smaller than for the standard one, in which the ROM prior is conditioned on uncorrected data. Clearly, using the estimated ROM error to correct the posterior yields a result close to the ground truth even though the ROM prior is used.
The corresponding predictive states for $m=5$ are visible in Figure 7. The ROM error estimate was computed with $12$ training points evenly spaced throughout the domain. In the result for the standard approach, i.e. the yellow curve $\tilde{\bm{u}}_{Y}$, the ROM prior is scaled with $\rho$ to account for the data as well as possible. The estimated model error is quite large as the ROM approximation error is part of it. Hence, the resulting variance is also relatively large compared to the prior and the full order predictive posterior. The statROM approach with the adjoint ROM error estimator, i.e. the purple line $\tilde{\bm{u}}_{Y;d_{r}}$, results in a smaller model error and hence, a much smaller predictive posterior variance with a mean much closer to the ground truth.
@@
One may observe that the proposed statROM approach yields more accurate results. The key insight of the results in Figure 7 is the data correction induced by the statROM update. Clearly, the yellow curve (statFEM) is the ROM prior scaled by $\rho$ so that it passes through the measurement data. Because of the ROM-corrected data model, here the resulting marginal likelihood (67), introduces a bias as large as the ROM error at the data points: the hyperparameters are estimated in order to comply with the corrected data, i.e. data close to the ROM prior. Therefore, the estimated model error GP ($\bm{d}$) of $\tilde{\bm{u}}_{Y_{d_{r}}}$ exhibits a very small $\sigma_{d}$: The ROM-prior lies close to the corrected data. The variance for the corrected posterior, as visible in Fig. 7, is rather small.
@@
On the other hand, the estimated model error GP $\bm{d}$ of $\tilde{\bm{u}}_{Y}$, i.e. for the uncorrected posterior, exhibits a rather large $\sigma_{d}$: The (ROM)-prior doesn’t lie on the actual measured data. This reflects a large posterior variance as visible in Fig. 7.